OpenStudy (rock_mit182):

Please someone help to figure this out: the problem says that the force between two spheres is 4N they're separeted by 10 cm. If the distance is two times less the initial what will happen with the magnitud of the force between them. columb's laws = F =k\frac{ q*q }{ r^2 }

OpenStudy (anonymous):

i can help you

OpenStudy (rock_mit182):

\[4 = k \frac{ q*q }{r^2 } \]

OpenStudy (rock_mit182):

0,1 meters for the initial values

OpenStudy (rock_mit182):

now if the radius is multiplied by 1/2

OpenStudy (anonymous):

yes

OpenStudy (rock_mit182):

the radius would be 1/20, right ?

OpenStudy (anonymous):

yea you got it keep going

OpenStudy (rock_mit182):

im stuck

OpenStudy (rock_mit182):

i think the right options are 8N or 16 N

OpenStudy (anonymous):

16 N

OpenStudy (rock_mit182):

because r is squared, right ?

OpenStudy (anonymous):

yea

OpenStudy (rock_mit182):

now if de charge of the two spheres is two times each one, mantaining the initial radius an force, the answer would be the same ?

OpenStudy (rock_mit182):

\[4 = k \frac{ 2q*2q }{ (0,1)^2 }\]

OpenStudy (anonymous):

okay so the initial force between the two charges separated by a distance d is F = 2kQ2 / d2 . After touching, the charges become Q/2 and 5Q/4 and the force is 5kQ2 / 8d2 = 5F / 16 . so the answer would be 5F / 16

OpenStudy (anonymous):

you get it?

OpenStudy (rock_mit182):

what did you do ?

OpenStudy (rock_mit182):

did you see my last equation ?

OpenStudy (rock_mit182):

i curious about what just did happen

OpenStudy (anonymous):

which one?

OpenStudy (rock_mit182):

\[4 =k \frac{ 2q *2q }{ (0,1)^2 }\]

OpenStudy (rock_mit182):

F1 =2 so F2 = ? ; keeping in mind that Im using the initial values but double charge ( each one)

OpenStudy (rock_mit182):

im sorry F1 =4 N

OpenStudy (anonymous):

oka what is the question asking? id dont see

OpenStudy (rock_mit182):

F2 is incresing respect to one by a factor of ?

OpenStudy (anonymous):

is it multiple choice i think its increasing by 2 but im not sure you might wanna ask someone else for a second opinion

OpenStudy (rock_mit182):

and i think is by a factor of 4

OpenStudy (rock_mit182):

anyway thanks a lot for helping me out

OpenStudy (anonymous):

no problem anytime

OpenStudy (rock_mit182):

@IrishBoy123 any thoughts ?

OpenStudy (irishboy123):

you have an inverse square relationship here so you can say \(\large F = \frac{C}{r^2} \) where C is a constant ie, \(\large C = F \ r^2\) so, applying this: \(\large C = 4 \times 10^2 = F \times 5^2\) what is F? [PS "two times less the" doesn't mean very much so i assumed you meant distance is halved]

OpenStudy (rock_mit182):

yeah that is what i meant

OpenStudy (rock_mit182):

distance is halved

OpenStudy (irishboy123):

so \[4 \times 10^2 = ?? \times 5^2\]